# Topological consequences of null-geodesic refocusing and applications to $Z^{x}$ manifolds

### v4

@@ 2 Main Results
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with an analytic metric $h$, then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
@@ 4 Observer-Refocusing Spacetimes
4.33 Remark.
In general, the spacetimes $(X,g)$ we consider will not be null-geodesically complete, so expressions of the type $\alpha_{v}(t)$ might become undefined for large $t$. However, every globally hyperbolic spacetime $(X,g)$ is conformally equivalent to a spacetime $(X^{\prime},g^{\prime})$ which is null-geodesically complete (see [17], where this is proven for all strongly causal spacetimes). Our proofs rely only on the causal structure and topology of our spacetime, both of which are preserved by conformal isomorphisms. So, when convenient, we will assume without loss of generality that our spacetime $(X,g)$ is null-geodesically complete.
@@
Let $(X,g)$ be a globally hyperbolic spacetime, $p\in X$, let $L$ be a lightsphere at $p$, let $A\subseteq L$ be a dense subset of $L$, and let $T>0$. If for every null-geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in A$ we have that $\alpha(T)\in I^{+}(p)$, then $X$ has a compact Cauchy surface.
@@
By Remark 4.33, we can assume that $(X,g)$ is null-geodesically complete. Let $v\in L$ and let $v_{n}$ be a sequence in $A$ converging towards $v$. Let $t_{n}$ be such that for every $n\in\mathbb{N}$ we have that $\alpha_{v_{n}}(t_{n})$ is the null-cut point of $\alpha_{v_{n}}|_{[0,t_{n}]}$. By assumption the sequence $(t_{n})_{n\in\mathbb{N}}$ is bounded by $T$. Without loss of generality assume that the sequence converges to some value $t$, else go to a subsequence. Since the null-cut locus $C_{N}^{+}(p)$ is closed (see [2, Proposition 6.5]), we have $\alpha_{v}(t)\in C_{N}^{+}(p)$. Therefore $\alpha_{v}(t^{\prime})\in I^{+}(p)$ for all $t^{\prime}>t$. The result follows by applying Lemma 4.30. ∎
4.35 Lemma.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$ and let $T>0$. If $L$ can be written as a countable union of closed sets
$L=\bigcup_{i\in I}K_{i}$
such that for each $i\in I$ and each $v\in\mathrm{int}(K_{i})$ we have $\alpha_{v}(T)\in I^{+}(p)$, then the Cauchy surface of $(X,g)$ is compact.
Proof.
By Remark 4.33, we can assume that $(X,g)$ is null-geodesically complete. Each set $K_{i}$ is closed in $L$, hence the boundaries $\partial_{L}(K_{i})$ are nowhere dense in $L$. Let $A=L\setminus\bigcup_{i\in I}\partial_{L}(K_{i})$, then $A$ is the complement of a countable union of nowhere dense sets, hence $A$ is dense by the Baire category theorem. Further, if $v\in A$, then there is some $i\in I$ with $v\in K_{i}\setminus\partial_{L}(K_{i})=\mathrm{int}_{L}(K_{i})$, and hence $\alpha_{v}(t)\in I^{+}(p)$ for all $t\geq T$. Since $A$ is dense, $X$ has a compact Cauchy surface by Lemma 4.34.
@@
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$ and let $\gamma$ be some timelike curve in $X$. Let $L$ be a lightsphere at $p$ and $v\in L$ and let $T>0$. Let $A=\{v\in L\mid\alpha_{v}|_{(0,T]}\text{ intersects }\gamma\}$ and $B=\{v\in L\mid p\text{ has a conjugate point }q\in\mathrm{im}(\gamma)\text{ along }\alpha_{v}\})$. Then $\mathrm{int}(A)\subseteq B$.
@@
By Remark 4.33, we can assume that $(X,g)$ is null-geodesically complete. Let $v\in\mathrm{int}(A)$. Since $\dim(X)\geq 3$ implies $\dim(L)\geq 1$, there is some sequence $(v_{n})_{n\in\mathbb{N}}$ in $A$ converging to $v$. For each $n\in\mathbb{N}$ let $t_{n}\in(0,T])$ and $q_{n}\in\mathrm{im}(\gamma)$ be such that $\alpha_{v_{n}}(t_{n})=q_{n}$. Since all $q_{n}$ are contained in the compact set $\{\exp(tv)\ \mid v\in L,t\in[0,T]\}$, we can assume that $(t_{n})_{n\in\mathbb{N}}$ and $(q_{n})_{n\in\mathbb{N}}$ converge to some $t$ and $q$ respectively, else go to convergent subsequences. By Lemma 4.36 there exists a $\mathcal{C}^{1}$ curve $\beta:[0,1]\to\mathcal{N}_{p}X$ with $\beta(0)=tv$ and $\beta^{\prime}(0)\neq 0$ and such that for every $\varepsilon>0$, there are infinitely many $t_{n}v_{n}\in\beta([0,\varepsilon])$. Consider the $\mathcal{C}^{1}$ curve $\alpha:[0,1]\to X$ defined through $\alpha(s)\coloneqq\exp(\beta(s))$. Then for every $\varepsilon>0$, there are infinitely many intersections of $\alpha|_{[0,\varepsilon]}$ and $\gamma$. Let $t$ be such that $\alpha(0)=q=\gamma(t)$. Since both $\alpha$ and $\gamma$ are $\mathcal{C}^{1}$, the fact that $\alpha|_{[0,\varepsilon]}$ intersects $\gamma$ an infinite number of times for every $\varepsilon>0$ implies that $\alpha^{\prime}(0)$ and $\gamma^{\prime}(t)$ are linearly dependent. Since $\gamma^{\prime}(t)\neq 0$ there is some $\lambda\in\mathbb{R}$ with $\alpha^{\prime}(0)=\lambda\cdot\gamma^{\prime}(t)$. If $\lambda\neq 0$, this would be a contradiction to the fact that $\alpha^{\prime}(0)$ is not timelike due to Lemma 4.37. So $\lambda=0$ and hence $\alpha^{\prime}(0)=0$. But $\alpha^{\prime}(0)=d\exp_{tv}(\beta^{\prime}(0))$ Since $\beta^{\prime}(0)\neq 0$ by construction, we conclude that $d\exp_{tv}$ is singular. This means that $q=\exp(tv)$ is conjugate along $\alpha_{v}$, and hence $v\in B$. ∎
@@
Using similar arguments as in Lemma 4.38 compactness of $\mathrm{im}(\gamma)$ ensures that there is some lightsphere $L$ at $p$ and some $T>0$ such that for all $v\in L$ there is some $t\in(0,T]$ with $\alpha_{v}(t)\in\mathrm{im}(\gamma)$. Now we can use Theorem 4.40. ∎
@@
One can carefully check the proof of Lemma 4.39 and Theorem 4.40 to see that if we drop the requirement of a uniform bound $T>0$ from Theorem 4.40 there will still be an open dense subset $A\subset L$ such that all null-geodesics $\alpha$ starting at $p$ with initial velocity in $A$ eventually enter $I^{+}(p)$, but this is not enough to conclude that we have a compact Cauchy surface.
@@ 5 The Analytic Case
Let $(X,g)$ be a globally hyperbolic spacetime. Let $p\in X$ and let $L$ be a lightsphere at $p$. Let $v\in L$ such that the geodesic $\alpha_{v}$ has at least $k$ points conjugate to $p$. Then there is some neighborhood $U\subseteq L$ of $v$ such that for each $w\in U$, the null-geodesic $\alpha_{w}$ has at least $k$ conjugate points (counted with multiplicity). Let $\lambda_{1}(w)\leq\lambda_{2}(w)\leq\dots\leq\lambda_{k}(w)$ be the parameter values at which $\alpha_{w}$ hits a conjugate point. Then the functions $\lambda_{j}$ are continuous on $U$ and they are smooth on an open, dense subset of $U$.
@@
The existence of the neighborhood $U$ and the continuity of the functions $\lambda_{j}$ are essentially the content of [31, Proposition 4.2]. The fact that each $\lambda_{j}$ is smooth on an open dense set of $U$ follows from [31, Theorem 5.3]: The set $\{\lambda_{j}(v)v\mid v\in U\}$ is a subset of the $T$-conjugate locus as defined in [31]. The regular null $T$-conjugate locus is an open dense subset in the null $T$-conjugate locus which is a smooth submanifold of $\mathcal{N}_{p}X$ transverse to rays in $\mathcal{N}_{p}X$. This implies that for every $j\in\{1,\dots,l\}$ there is an open dense subset $U_{j}\subset U$ such that $\{\lambda_{j}(v)v\mid v\in U_{j}\}$ is a smooth submanifold of $\mathcal{N}_{p}X$ transverse to the rays in $\mathcal{N}_{p}$. By the implicit function theorem the map $v\mapsto\lambda_{j}(v)$ is smooth on $U_{j}$. ∎
@@
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$. Let $\gamma$ be some timelike curve in $X$, let $p\in X$ and let $L$ be a lightsphere at $p$. Let $U\subseteq L$ be some open set. Let $f:U\to\mathbb{R}^{+}$ be some lower semi-continuous function such that for all $v\in U$ we have that $\exp(f(v)v)\in\mathrm{im}(\gamma)$ is a point conjugate to p along the geodesic $\alpha_{v}$. Then there is some non-empty open set $V\subseteq U$ on which $\exp(f(v)v)$ is constant.
@@
Let the functions $\lambda_{j}$ be as in Theorem 5.48. For $v\in U$ we have that $\exp(f(v)v)$ is a conjugate point along $\alpha_{v}$. Let $k(v)$ denote the minimal $k$ with $\lambda_{k}(v)=f(v)$. We will show that $v\mapsto k(v)$ is lower semi-continuous. Let $v_{n}$ be some sequence in $U$ converging to some $v\in U$ and let $N\in\mathbb{N}$ with $k_{n}\coloneqq k(v_{n})\leq N$ for all $n\in\mathbb{N}$. Then $f(v_{n})=\lambda_{k_{n}}(v_{n})\leq\lambda_{N}(v_{n})$ for all $n\in\mathbb{N}$. By continuity of $\lambda_{N}$ and lower semi-continuity of $f$, we get that $\lambda_{k(v)}(v)=f(v)\leq\lambda_{N}(v)$. Since $k(v)$ is the smallest $k$ with $\lambda_{k}(v)=f(v)$, we must have $k(v)\leq N$. Therefore the function $v\mapsto k(v)$ is lower semi-continuous. For $i\in\mathbb{N}$, let $A_{i}=\{v\in U\mid k(v)\leq i\}$ and note that $U=\bigcup_{i\in\mathbb{N}}A_{i}$. By the Baire category theorem, there is some $i\in\mathbb{N}$ such that $A_{i}$ has non-empty interior. Take $i\in\mathbb{N}$ minimal with this property. Then $V\coloneqq\mathrm{int}(A_{i})\setminus A_{i-1}$ is open. If $\mathrm{int}(A_{i})\setminus A_{i-1}=\emptyset$, then we would have had $\mathrm{int}(A_{i})\subseteq A_{i-1}$ and hence $\mathrm{int}(A_{i})\subseteq\mathrm{int}(A_{i-1})$, contradicting the minimality of $i$. So $V$ is non-empty. For $v\in V$ we have that $f(v)=\lambda_{i}(v)$ and by Theorem 5.48 we can shrink $V$ to some smaller non-empty open set so that $f|_{V}=\lambda_{i}|_{V}$ is smooth. Assume that this $V$ is connected else take a connected component of it instead. Let $s\mapsto v_{s}$ be any $\mathcal{C}^{1}$ curve in $V$ and let us denote $v=f(v_{0})v_{0}$ and $w=\frac{d}{ds}|_{s=0}(f(v_{s})v_{s})$. Then, by Lemma 4.37, $\frac{d}{ds}|_{s=0}\exp(f(v_{s})v_{s})=d\exp_{v}(w)$ is not timelike. But by construction $f(v_{s})\in\mathrm{im}(\gamma)$ for all $s$, so $\frac{d}{ds}f(v_{s})$ is some scalar multiple of a timelike vector. This is only possible if $\frac{d}{ds}|_{s=0}\exp(f(v_{s})v_{s})=d\exp_{v}(w)=0$. Since on $V$ all directional derivatives of $v\mapsto\exp(f(v)v)$ vanish, we conclude that $v\mapsto\exp(f(v)v)$ is constant on $V$. ∎
@@
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$ which is observer-refocusing with respect to a point $p\in X$ and a timelike curve $\gamma$ in $X$. Let $L$ be a lightsphere at $p$. Then there exists a non-empty open subset $V\subseteq L$ and a point $q\in\mathrm{im}(\gamma)$ such that for all $v\in V$ there is some $t>0$ with $\alpha_{v}(t)=q$
@@
As shown in the proof of Theorem 4.40, there is an open, dense subset $L^{\prime}\subseteq L$ such that for all $v\in L^{\prime}$, $p$ has some conjugate point $q$ in the intersection of $\alpha_{v}$ and $\gamma$. We define $f:L^{\prime}\to\mathbb{R}^{+}$ through $\exp(f(v)v)=\text{first conjugate point }q\in\mathrm{im}(\gamma)\text{ along }\alpha_{v}.$ It is easy to see $f$ is lower semi-continuous. By Lemma 5.49, $v\mapsto\exp(f(v)v)$ is constant on some non-empty open set $V\subseteq L^{\prime}$. In particular, there is some $q\in\mathrm{im}(\gamma)$ such that for all $v\in V$ there is some $t=f(v)>0$ such that $\alpha_{v}(t)=\exp(tv)=q$. ∎
@@
By [33, Theorem 2.10], there exists an analytic temporal function $\mathcal{T}:X\to\mathbb{R}$ whose level sets are Cauchy surfaces. Let $p\in X$ and let $\gamma$ be a timelike curve in $X$ so that $(X,g)$ is observer-refocusing with respect to $p$ and $\gamma$. Let $L\subset T_{p}X$ be some lightsphere at $p$ which is also an analytic submanifold of $T_{p}X$, i.e. a lightsphere such that the inclusion $\iota:L\to T_{p}X$ is analytic. Let $D\subset L\times\mathbb{R}$ be the maximal domain of the function $(v,t)\mapsto\exp_{p}(tv)$. Note that the map $H:D\to T_{p}X$ defined through $H(v,t)=t\cdot\iota(v)$ is analytic. Consider the map
@@
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with an analytic metric $h$, then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
@@
In particular, any $Z^{x}$ manifold with analytic metric is a $Y^{x}_{l}$ manifold for some $l>0$ (the $1$-dimensional case is trivially true even without analyticity).

### v5

@@ 2 Main Results
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with an analytic metric $h$. Then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
@@ 4 Observer-Refocusing Spacetimes
Let $(X,g)$ be a globally hyperbolic spacetime, $p\in X$, let $L$ be a lightsphere at $p$, let $L^{\prime}\subseteq L$ be a dense subset of $L$, and let $T>0$ so that $\alpha_{v}(T)$ is defined for all $v\in L$. If for every null-geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in A$ we have that $\alpha(t)\in I^{+}(p)$ for all $t>T$, then $X$ has a compact Cauchy surface.
@@
Let $v\in L$ and let $v_{n}$ be a sequence in $L^{\prime}$ converging towards $v$. Let $t_{n}$ be such that for every $n\in\mathbb{N}$ we have that $\alpha_{v_{n}}(t_{n})$ is the null-cut point of $\alpha_{v_{n}}|_{[0,t_{n}]}$. By assumption the sequence $(t_{n})_{n\in\mathbb{N}}$ is bounded by $T$. Without loss of generality assume that the sequence converges to some value $t$, else go to a subsequence. Since the null-cut locus $C_{N}^{+}(p)$ is closed (see [2, Proposition 6.5]), we have $\alpha_{v}(t)\in C_{N}^{+}(p)$. Therefore $\alpha_{v}(t^{\prime})\in I^{+}(p)$ for all $t^{\prime}>t$. The result follows by applying Lemma 4.30. ∎
@@
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$ and let $\gamma$ be some timelike curve in $X$. Let $L$ be a lightsphere at $p$ and let $T>0$ be such that $\alpha_{v}(T)$ is defined for all $v\in L$. Let $A=\{v\in L\mid\alpha_{v}|_{(0,T]}\text{ intersects }\gamma\}$ and $B=\{v\in L\mid p\text{ has a conjugate point }q\in\mathrm{im}(\gamma)\text{ along }\alpha_{v}|_{(0,T]}\}$. Then $\mathrm{int}(A)\subseteq B$.
@@
Let $v\in\mathrm{int}(A)$. Since $\dim(X)\geq 3$ implies $\dim(L)\geq 1$, there is some sequence $(v_{n})_{n\in\mathbb{N}}$ in $A$ converging to $v$. For each $n\in\mathbb{N}$ let $t_{n}\in(0,T])$ and $q_{n}\in\mathrm{im}(\gamma)$ be such that $\alpha_{v_{n}}(t_{n})=q_{n}$. Since all $q_{n}$ are contained in the compact set $\{\exp(tv)\ \mid v\in L,t\in[0,T]\}$, we can assume that $(t_{n})_{n\in\mathbb{N}}$ and $(q_{n})_{n\in\mathbb{N}}$ converge to some $t$ and $q$ respectively, else go to convergent subsequences. By Lemma 4.34 there exists a $\mathcal{C}^{1}$ curve $\beta:[0,1]\to\mathcal{N}_{p}X$ with $\beta(0)=tv$ and $\beta^{\prime}(0)\neq 0$ and such that for every $\varepsilon>0$, there are infinitely many $t_{n}v_{n}\in\beta([0,\varepsilon])$. Consider the $\mathcal{C}^{1}$ curve $\alpha:[0,1]\to X$ defined through $\alpha(s)\coloneqq\exp(\beta(s))$. Then for every $\varepsilon>0$, there are infinitely many intersections of $\alpha|_{[0,\varepsilon]}$ and $\gamma$. Let $t$ be such that $\alpha(0)=q=\gamma(t)$. Since both $\alpha$ and $\gamma$ are $\mathcal{C}^{1}$, the fact that $\alpha|_{[0,\varepsilon]}$ intersects $\gamma$ an infinite number of times for every $\varepsilon>0$ implies that $\alpha^{\prime}(0)$ and $\gamma^{\prime}(t)$ are linearly dependent. Since $\gamma^{\prime}(t)\neq 0$ there is some $\lambda\in\mathbb{R}$ with $\alpha^{\prime}(0)=\lambda\cdot\gamma^{\prime}(t)$. If $\lambda\neq 0$, this would be a contradiction to the fact that $\alpha^{\prime}(0)$ is not timelike due to Lemma 4.35. So $\lambda=0$ and hence $\alpha^{\prime}(0)=0$. But $\alpha^{\prime}(0)=d\exp_{tv}(\beta^{\prime}(0))$ Since $\beta^{\prime}(0)\neq 0$ by construction, we conclude that $d\exp_{tv}$ is singular. This means that $q=\exp(tv)$ is conjugate along $\alpha_{v}$, and hence $v\in B$. ∎
4.38 Remark.
In general, the spacetimes $(X,g)$ we consider will not be null-geodesically complete, so expressions of the type $\alpha_{v}(t)$ might become undefined for large $t$. However, every globally hyperbolic spacetime $(X,g)$ is conformally equivalent to a spacetime $(X^{\prime},g^{\prime})$ which is null-geodesically complete (see [17], where this is proven for all strongly causal spacetimes). Conformal isomorphisms preserve both the causal structure and topology of our spacetime. Therefore, when appropriate, we will assume without loss of generality that our spacetime $(X,g)$ is null-geodesically complete.
@@
By Remark 4.38, we can assume that $(X,g)$ is null-geodesically complete. Using similar arguments as in Lemma 4.36 compactness of $\mathrm{im}(\gamma)$ ensures that there is some lightsphere $L$ at $p$ and some $T>0$ such that for all $v\in L$ there is some $t\in(0,T]$ with $\alpha_{v}(t)\in\mathrm{im}(\gamma)$. Now we can use Theorem 4.39. ∎
@@
One can carefully check the proof of Lemma 4.37 and Theorem 4.39 to see that if we drop the requirement of a uniform bound $T>0$ from Theorem 4.39 there will still be an open dense subset $L^{\prime}\subset L$ such that all null-geodesics $\alpha_{v}$ starting at $p$ with initial velocity $v\in L^{\prime}$ have a conjugate point along one of the curves $\gamma_{i}$, hence eventually enter $I^{+}(p)$. In the proof of Theorem 4.39 one can define the countable family $(A_{(i,N)})_{i\in I,N\in\mathbb{N}}$ through $A_{(i,N)}=\{v\in L\mid\alpha_{v}|_{(0,N]}\text{ intersects }\gamma_{i}\}$. Then the union $L^{\prime}$ of the interiors of all the $A_{(i,N)}$ is open and dense in $L$. But this is not enough to conclude that we have a compact Cauchy surface since we cannot apply 4.33 without a uniform bound.
@@ 5 The Analytic Case
Let $(X,g)$ be a globally hyperbolic spacetime. Let $p\in X$ and let $L$ be a lightsphere at $p$. Let $v\in L$ such that the geodesic $\alpha_{v}$ has at least $k$ points conjugate to $p$. Then there is some neighborhood $U\subseteq L$ of $v$ such that for each $w\in U$, the null-geodesic $\alpha_{w}$ has at least $k$ conjugate points (counted with multiplicity). Let $0<\lambda_{1}(w)\leq\lambda_{2}(w)\leq\dots\leq\lambda_{k}(w)$ be the parameter values at which $\alpha_{w}$ hits a conjugate point. Then the functions $\lambda_{j}$ are continuous on $U$ and they are smooth on an open, dense subset of $U$.
@@
The existence of the neighborhood $U$ and the continuity of the functions $\lambda_{j}$ are essentially the content of [30, Proposition 4.2]. The fact that each $\lambda_{j}$ is smooth on an open dense set of $U$ follows from [30, Theorem 4.3]: The set $\{\lambda_{j}(v)v\mid v\in U\}$ is a subset of the $T$-conjugate locus as defined in [30]. The regular null $T$-conjugate locus is an open dense subset in the null $T$-conjugate locus which is a smooth submanifold of $\mathcal{N}_{p}X$ transverse to rays in $\mathcal{N}_{p}X$. This implies that for every $j\in\{1,\dots,k\}$ there is an open dense subset $U_{j}\subseteq U$ such that $\{\lambda_{j}(v)v\mid v\in U_{j}\}$ is a smooth submanifold of $\mathcal{N}_{p}X$ transverse to the rays in $\mathcal{N}_{p}X$. By the implicit function theorem the map $v\mapsto\lambda_{j}(v)$ is smooth on $U_{j}$. ∎
@@
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$. Let $\gamma$ be some timelike curve in $X$, let $p\in X$ and let $L$ be a lightsphere at $p$. Let $U\subseteq L$ be some non-empty open set. Let $f:U\to\mathbb{R}^{+}$ be some lower semi-continuous function such that for all $v\in U$ we have that $\exp(f(v)v)\in\mathrm{im}(\gamma)$ is a point conjugate to $p$ along the geodesic $\alpha_{v}$. Then there is some non-empty open set $V\subseteq U$ on which $\exp(f(v)v)$ is constant.
@@
Let the functions $\lambda_{j}$ be as in Theorem 5.47. For $v\in U$ we have that $\exp(f(v)v)$ is a conjugate point along $\alpha_{v}$. Let $k(v)$ denote the minimal $k$ with $\lambda_{k}(v)=f(v)$. We will show that $v\mapsto k(v)$ is lower semi-continuous. Let $v_{n}$ be some sequence in $U$ converging to some $v\in U$ and let $N\in\mathbb{N}$ with $k_{n}\coloneqq k(v_{n})\leq N$ for all $n\in\mathbb{N}$. Then $f(v_{n})=\lambda_{k_{n}}(v_{n})\leq\lambda_{N}(v_{n})$ for all $n\in\mathbb{N}$. By continuity of $\lambda_{N}$ and lower semi-continuity of $f$, we get that $\lambda_{k(v)}(v)=f(v)\leq\lambda_{N}(v)$. Since $k(v)$ is the smallest $k$ with $\lambda_{k}(v)=f(v)$, we must have $k(v)\leq N$. Therefore the function $v\mapsto k(v)$ is lower semi-continuous. For $i\in\mathbb{N}$, let $A_{i}=\{v\in U\mid k(v)\leq i\}$ and note that $U=\bigcup_{i\in\mathbb{N}}A_{i}$. By the Baire category theorem, there is some $i\in\mathbb{N}$ such that $A_{i}$ has non-empty interior. Take $i\in\mathbb{N}$ minimal with this property. Then $V\coloneqq\mathrm{int}(A_{i})\setminus A_{i-1}$ is open. If $\mathrm{int}(A_{i})\setminus A_{i-1}=\emptyset$, then we would have had $\mathrm{int}(A_{i})\subseteq A_{i-1}$ and hence $\mathrm{int}(A_{i})\subseteq\mathrm{int}(A_{i-1})$, contradicting the minimality of $i$. So $V$ is non-empty. For $v\in V$ we have that $f(v)=\lambda_{i}(v)$ and by Theorem 5.47 we can shrink $V$ to some smaller non-empty open set so that $f|_{V}=\lambda_{i}|_{V}$ is smooth. Assume that this $V$ is connected else take a smaller $V$ that is instead. Let $s\mapsto v_{s}$ be any $\mathcal{C}^{1}$ curve in $V$ and let us denote $v=f(v_{0})v_{0}$ and $w=\frac{d}{ds}|_{s=0}(f(v_{s})v_{s})$. Then, by Lemma 4.35, $\frac{d}{ds}|_{s=0}\exp(f(v_{s})v_{s})=d\exp_{v}(w)$ is not timelike. But by construction $\exp(f(v_{s})v_{s})\in\mathrm{im}(\gamma)$ for all $s$, so $\frac{d}{ds}\exp(f(v_{s})v_{s})$ is some scalar multiple of a timelike vector. This is only possible if $\frac{d}{ds}|_{s=0}\exp(f(v_{s})v_{s})=d\exp_{v}(w)=0$. Since on $V$ all directional derivatives of $v\mapsto\exp(f(v)v)$ vanish, we conclude that $v\mapsto\exp(f(v)v)$ is constant on $V$. ∎
@@
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$ which is observer-refocusing with respect to a point $p\in X$ and a timelike curve $\gamma$ in $X$. Let $L$ be a lightsphere at $p$. Then there exists a non-empty open subset $V\subseteq L$ and a point $q\in\mathrm{im}(\gamma)$ such that for all $v\in V$ there is some $t>0$ with $\alpha_{v}(t)=q$.
@@
As discussed in Remark 4.45, there is an open, dense subset $L^{\prime}\subseteq L$ such that for all $v\in L^{\prime}$, $p$ has some conjugate point $q$ in the intersection of $\alpha_{v}$ and $\gamma$. We define $f:L^{\prime}\to\mathbb{R}^{+}$ through $\exp(f(v)v)=\text{first conjugate point }q\in\mathrm{im}(\gamma)\text{ along }\alpha_{v}.$ It is easy to see that $f$ is lower semi-continuous. By Lemma 5.48, $v\mapsto\exp(f(v)v)$ is constant on some non-empty open set $V\subseteq L^{\prime}$. In particular, there is some $q\in\mathrm{im}(\gamma)$ such that for all $v\in V$ there is some $t=f(v)>0$ such that $\alpha_{v}(t)=\exp(tv)=q$. ∎
@@
By [32, Theorem 2.10], there exists an analytic temporal function $\mathcal{T}:X\to\mathbb{R}$ whose level sets are Cauchy surfaces. Let $p\in X$ and let $\gamma$ be a timelike curve in $X$ so that $(X,g)$ is observer-refocusing with respect to $p$ and $\gamma$. Let $L\subset T_{p}X$ be some lightsphere at $p$ which is also an analytic submanifold of $T_{p}X$, i.e. a lightsphere such that the inclusion $\iota:L\to T_{p}X$ is analytic. Let $D\subset L\times\mathbb{R}$ be the maximal domain of the function $(v,t)\mapsto\exp_{p}(tv)$. Note that the map $H:D\to T_{p}X$ defined through $H(v,t)=t\cdot\iota(v)$ is analytic. Consider the map
@@
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with an analytic metric $h$. Then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
@@
In particular, any $Z^{x}$ manifold with analytic metric is a $Y^{x}_{l}$ manifold for some $l>0$ (the $1$-dimensional case is trivially true even without analyticity). Since Corollary 5.52 might be of independent interest, a self-contained proof of this result, avoiding Lorentzian geometry entirely, is supplied in the appendix.